OpenStudy (anonymous):

if the remainder on division of x^3+2x^2+kx+3 by x-3 is 21,find the quotient and the value of k. Hence,find the zeroes of the cubic polynomial x^3+2x^2+kx-18.

OpenStudy (shubhamsrg):

it means that if f(x) = x^3+2x^2+kx+3 by x-3 then f(3) = 21 (since x-3 leaves remainder 21 on division) make that substitution and solve for k. then you use long division up next to solve for quotient ,,and then the roots.

OpenStudy (anonymous):

\[x^3+2x^2+kx+3=(x-3)(x^2+ax+b)+21 \\ u \ can \ utilize \ this\]

OpenStudy (anonymous):

Oh gosh...this will take a while... First, set up the synthetic division. 3| 1 2 k 3 3 15 3(k + 15) 1 5 (k + 15) 21 Therefore, set up an equation. 3(k + 15) + 3 = 21 3(k + 15) = 18 k + 15 = 6 k = -9 x^3 + 2x^2 - 9x + 3

OpenStudy (anonymous):

Now that you have the value of k, use the rational root theorem and Descartes Rule of Signs to decrease the degree to a quadratic.